OpenStudy (anonymous):

What is the common ratio of this geometric sequence? 1 80 2 40 3 20 4 10 5 5 A. -40 B. 0.5 C. 2 D. 40

jimthompson5910 (jim_thompson5910):

common ratio = (second term)/(first term) = 40/80 = 1/2 = 0.5 this means you multiply each term by 0.5 to get the next term

jimthompson5910 (jim_thompson5910):

80 times 0.5 = 40 40 times 0.5 = 20 etc etc